Michael invested his savings in two investment funds. The amount he invested in Fund A was $3000
less than the amount he invested in Fund B. Fund A returned a 8%
profit and Fund B returned a 6%
profit. How much did he invest in Fund B, if the total profit from the two funds together was $1160?

Answer by mananth(16946)

6.00% --------- x
8.00% --------- x + 3000

Total invested = (x+x+ 3000 )
Interest on total investment = 1160

6.00% x+ 8.00% (+x 3000 )= 1160

multiply by 100
we get
6 x+ 8 ( 1 x + 3000 )= 116000
6 x+ 8 x+ 24000 = 116000

14 x= 116000 -24000
x= 92000 / 14
x= 6571.43

$ 6571.43 @ 6.00%

1 x+ 3000 = 9571.43

$ 9571.43 @ 8.00%

CHECK
6571.43 * 6.00% + 9571.43 * 8.00%
( 394.29 + 765.71 = 1160